OpenStudy (anonymous):

sove for x 12-3x-x=8

OpenStudy (anonymous):

12-3x-x=8 12-4x=8 -4x=-4 x=1

OpenStudy (anonymous):

The answer is x=1

OpenStudy (anonymous):

@RavenWinchester beat me to it

OpenStudy (anonymous):

Haha sorry. Nice to meet ya, though. :) Have a great night.

OpenStudy (anonymous):

It's okay but I did my equation a different way.

OpenStudy (anonymous):

How'd you do yours?

OpenStudy (anonymous):

12-3x-x=8 12-3=9 9*1=9-1=8 I tried to find out what number was more close to 8 when you multiplied a number by 9

OpenStudy (anonymous):

Creative I like it :)

OpenStudy (anonymous):

ty